OpenStudy (anonymous):

DIFFERENTIAL EQUATIONS Find the function f s.t. f'(x) = f(x)*(1- f(x)) and f(0) = 1/2. Does this want me to solve for y, if I say y = f(x)? I show you the eqn I used:

OpenStudy (anonymous):

@Mertsj @Zarkon @dpaInc @jim_thompson5910 @lgbasallote @KingGeorge @AccessDenied @SmoothMath @eliassaab @asnaseer @.Sam. @kropot72 @eigenschmeigen @satellite73 @amistre64 @TuringTest @shivam_bhalla @experimentX @FoolForMath @Callisto @phi @mahmit2012 @Chlorophyll @ParthKohli @saifoo.khan @robtobey @radar

OpenStudy (anonymous):

\[\frac{dy}{dx} = y(1-y)\] Where y = f(x)

OpenStudy (anonymous):

@saifoo.khan are you experienced with this?

OpenStudy (anonymous):

This wants me to solve for y =, right?

OpenStudy (anonymous):

But here is the problem: \[\ln|1 - y| + \ln|y| = x + \ln(1/4)\] How the heck can you solve for y there?

OpenStudy (anonymous):

Introduce ln to every term then differentiate. Derivative of \[\ln x= 1/x\] But then we have y. So remember to multiply with y' or dy/dx then solve for y' .Need me solve it?

OpenStudy (anonymous):

The absolute doesn't make a difference because ln x has to be greater than zero

OpenStudy (anonymous):

If it says find the function f (or I said f(x) = y), does that not mean I need y =?

OpenStudy (saifoo.khan):

Sorry no.

OpenStudy (anonymous):

Hold on, So what is the function you wrote down equal to?

sam (.sam.):

you have to find f(x), \[f'(x) = f(x)(1- f(x)) ~and ~f(0) = 1/2\] \[\frac{dy}{dx}=y(1-y)\] \[\begin{array}{l} \text{rearrange} \\ \frac{dy}{dx}\text{ = }-(y-1) y \\ \text{divide both sides by }-(y-1) y \\ -\frac{1}{(y-1) y}\frac{dy}{dx}\text{ = }1 \\ \text{integrate both sides } \\ \int\limits -\frac{1}{(y-1) y} \frac{dy}{dx}\, dx\text{ = }\int\limits 1 \, dx \\\end{array}\] you'll get \[-\ln (-y+1)+\ln (y)\text{ = }x+c\]

OpenStudy (anonymous):

WTH? Where did you get that (-) from?

sam (.sam.):

you rearrange it

OpenStudy (anonymous):

Ok I see it

OpenStudy (anonymous):

Oh so you use the quotient rule for logs afterwards to get rid of the possibility of there being a y^2 term

OpenStudy (anonymous):

U substitution @QRAwarrior That's what you use to integrate such a function.

sam (.sam.):

\[-\ln (-y+1)+\ln (y)\text{ = }x+c\] \[\ln(\frac{y}{1-y})=x+c\] \[\frac{y}{1-y}=e^{x+c}\]

OpenStudy (anonymous):

Wait, so then you do NOT get a "y =" expression, because of that (y)/(1 -y)..

sam (.sam.):

\[\frac{y}{1-y}=e^{x+c}\] \[y=e^{x+c}(1-y)\] \[y=e^{x+c}-ye^{x+c}\] \[y+ye^{x+c}=e^{x+c}\] \[y(1+e^{x+c}=e^{x+c}\] \[y=\frac{e^{x+c}}{1+e^{x+c}}\]

OpenStudy (anonymous):

Ok I do not want to look at that to its entirety until I try your rearranging strategem. Thanks for your help.

sam (.sam.):

yeah you should try rearranging them, then sub the f(0) = 1/2 and get c, then present the function with y and x